Solve the initial value problem below using the method of Laplace transforms. y - y = 4t - 10 e + y(0)= 0, y(O) = 13 Click
Answer #1

е ut - loet / 90)= 0, 90) — 13 4 ( t – ме?) 10 — 82 | у - y = че - Apply Laplace bls (9 – 9) я* (9) -я 90-90 - 219) = (x-1)